OpenStudy (edaiiiii):

If a piece of ribbon is 4 and 3 over 4. feet long, how long is it in inches?

OpenStudy (mathstudent55):



OpenStudy (mathstudent55):

Same method as all your previous problems. Multiply both sides by the same number.

OpenStudy (edaiiiii):

so i multiply 4 x 4 then 4 x 3

OpenStudy (skullpatrol):

Nope, you want to convert 4 3/4 ft into inches, right?

OpenStudy (edaiiiii):

yes

OpenStudy (edaiiiii):

how do i do that though

OpenStudy (skullpatrol):

4 3/4 ft = 19/4 ft X 12 in / 1 ft

OpenStudy (sila1453):

4 + 3/4 = 16/4 + 3/4 ya get 19/4 19/4 *12 = 228/4 = 57

OpenStudy (mathstudent55):

4 3/4 is a mixed numeral. Since 3/4 means three-fourths, 4 3/4 means 4 and 3 fourths. You can write it as a decimal number as 4.75

OpenStudy (mathstudent55):

You know 1 ft equals 12 inches. You want to know what 4.75 ft is equal to in inches, so multiply the conversion equation by 4.75 on both sides.

OpenStudy (mathstudent55):